OpenStudy (anonymous):

The Dinner Shoppe spends $23,850 monthly on supplies. If supplies represent 33 1/3% of total expenses, what are the annual total expenses to the nearest dollar?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Let \(x\) = the annual total expenses Setup a proportion and solve \(x\) \[\large \dfrac{33\frac{1}{3}}{100}=\frac{23850}{x}\]

OpenStudy (anonymous):

whats th e 1/3 stand for

ganeshie8 (ganeshie8):

\(33 \frac{1}{3}\) is mixed fraction you may read it ias 33 and a one third

ganeshie8 (ganeshie8):

In your calculations, you may replace it like this\[33\frac{1}{3}=33+\frac{1}{3}\]

OpenStudy (anonymous):

23850 = .333 X Expenses 71550 = Expenses (multiply both sides by 3 or divide by .333

OpenStudy (anonymous):

i dont know how to plug ur problem on my calculator

ganeshie8 (ganeshie8):

Looks good!

ganeshie8 (ganeshie8):

your method is fine

OpenStudy (anonymous):

it says its wrong i put it in as $ 71,550 says its wrong

ganeshie8 (ganeshie8):

just put "71550" without commas or dollar symbol

OpenStudy (anonymous):

still wrong hmm is it rounded to neearest dollar

ganeshie8 (ganeshie8):

take a screenshot and attach

OpenStudy (anonymous):

mm i cant im on the computer

OpenStudy (anonymous):

sw The Dinner Shoppe spends $23,850 monthly on supplies. If supplies represent 33 1/3% of total expenses, what are the annual total expenses to the nearest dollar?

ganeshie8 (ganeshie8):

then im not so sure, must be a grader problem

OpenStudy (anonymous):

2)Joint-Serve Prosthetics spends 43% of its total expenses for payroll. It increased its total expenses when it spent $250,000 for robotic assemblers. This changed the allocation for payroll to 38% of total expenses. What were its total expenses after purchasing the robotic assemblers? Hint: First find the total expense before the purchase of robotic assemblers.

OpenStudy (anonymous):

b payroll = .43 X Expenses payroll = .38 X (Expenses + 250000) Mult-ply out the second line Payroll = .38 X Expenses + 95000 subtract this from line 1 0 = .05 X Expenses - 95000 95000 = .05 Expenses Divide by .05 1,900,000 = Expenses Verify in equations Payroll = .43 X Expenses = .43 X 1,900,000 = $817,000 Payroll = .38 X (Expenses + 250000) = .38 X (1,900,000+250,000) = .38 X 2,150,000 = $817,000

OpenStudy (anonymous):

i puty it in as that $817,000 wrong too